# DialectBench: A NLP Benchmark for Dialects, Varieties, and Closely-Related Languages

### v1

@@ 1 Introduction
Figure 2: DialectBench language clusters with their variety counts per task. "Other" encompasses 18 clusters (full cluster list in Appendix Table 8).
@@ 2 DialectBench
Category
Task
Metric
Source Dataset
Structured
DEP parsing
UAS
Universal Dependency Zeman et al. (2021), TwitterAAE Blodgett et al. (2018), Singlish Wang et al. (2017)
Prediction
POS tagging
F1
Universal Dependency Zeman et al. (2021), Singlish Wang et al. (2017), Noisy Dialects Blaschke et al. (2023)
NER
F1
Wikiann Pan et al. (2017); Rahimi et al. (2019), Norwegian NER Johansen (2019)
Classification
DId
F1
MADAR Bouamor et al. (2018), DMT Jauhiainen et al. (2019), Greek Sababa and Stassopoulou (2018), DSL-TL Zampieri et al. (2023), Swiss Germans Scherrer et al. (2019)
SA
F1
TSAC Medhaffar et al. (2017), TUNIZI Fourati et al. (2021), DzSentiA Abdelli et al. (2019), SaudiBank Alqahtani et al. (2022), MAC Garouani and Kharroubi (2022), ASTD Nabil et al. (2015), AJGT Alomari et al. (2017), OCLAR Al Omari et al. (2019)
TC
F1
SIB-200 Adelani et al. (2023)
NLI
F1
XNLI Conneau et al. (2018) translate-test
Question
MRC
F1
Belebele Bandarkar et al. (2023)
Answering
EQA
Span F1
SDQA Faisal et al. (2021)
Generation
MT
BLEU
CODET Alam et al. (2023), TIL-MT Mirzakhalov (2021)
@@ 3.2 Training and Evaluation
We use six general approaches for task-specific model training:
@@
    Task
    In-variety FT
    In-cluster FT
    Combined FT
    Zero-shot
    No ref-erence
    In-context learning
    DEP
    $\checkmark$
    $\checkmark$
    POS
    $\checkmark$
    $\checkmark$
    NER
    $\checkmark$
    $\checkmark$
    EQA
    $\checkmark$
    $\checkmark$
    $\checkmark$
    MRC
    $\checkmark$
    NLI
    $\checkmark$
    TC
    $\checkmark$
    $\checkmark$
    SA
    $\checkmark$
    $\checkmark$
    DI
    $\checkmark$
    MT
    $\checkmark$
    $\checkmark$
@@ 4 Results
Structured prediction DEP parsing 16 40 64.3 southwestern shifted romance/brazilian portuguese 94.4 tupi-guarani subgroup i.a/mbyá guaraní (a:brazil) 9.0
POS tagging 17 51 72.1 norwegian/norwegian bokmål (m:written) 98.7 tupi-guarani subgroup i.a/mbyá guaraní (a:brazil) 1.9
@@
TC 15 38 77.7 sinitic/classical-middle-modern sinitic (o:traditional) 89.8 kurdish/central kurdish 19.4
DId 6 49 67.0 sinitic/mandarin chinese (a:taiwan, o:simplified) 98.6 southwestern shifted romance/portuguese (m:written) 17.4
@@
Qusetion Answering MRC 4 11 40.9 anglic/english 53.4 sotho-tswana (s.30)/southern sotho 29.0
@@ 4.1 Maximum Obtainable Scores
Here we provide key findings from our evaluation on each task. A task-specific summary is reported in Table 3. Detailed results comprising all tasks, models, language clusters and varieties are reported in Tables 10 to 20 in Appendix E.
@@
We present visualizations for the task-specific maximum scores. We show the one for Dependency Parsing in Fig. 3, where we observe that low-resource varieties from language clusters such as Tupi-Guarani (indigenous South American cluster), Saami and Komi (low-resource Uralic language clusters) have the lowest performance compared to Standard English and other closely related Germanic clusters. These low-resource varieties are also not included in the pretraining stage of our base language models (eg. mBERT). Furthermore, this trend is evident across all three structured prediction tasks. On the other hand, high-resource Indo-European languages such as Portuguese, French, and Norwegian usually perform better.
@@
For DId and SA, we generally collate different datasets for each language cluster and therefore, report the comparative classification results together. As a result, the locality level (e.g. city/region/country) also varies across cluster. For example, we report city-level DId results for Arabic and High German but country-level results for Portuguese, Spanish and English. In the case of SA, we have region/country-level results for Arabic varieties. For TC and NLI, we have the same set of clusters and varieties. However, we only report the zero-shot transfer performance from Standard English for NLI using the newly created translate-test NLI dataset.
@@
We generally do not see large gaps in performance within varieties in each language cluster. In EQA zero-shot experiments, English and its varieties have the highest performance overall and Korean varieties score the lowest. Combined fine-tuning boosts performance on all language clusters except in English. For MRC the performance threshold peaks at 53.4 for Standard English and the minimum we get is 29 for Southern Sotho. More detailed results are presented in Tables 19 and 15.
@@ 4.2 Dialectal Gap Across Language Clusters
Task
Gap metric Avg. val cluster (max) cluster (min)
@@ 5 Discussion
The highest-performing varieties are mostly standard high-resource languages and a few high-resource dialects (Norwegian dialects) whereas, the majority of the lowest-performing language variants are low-resourced varieties. This clear distinction of language varieties points towards the large existence of in-cluster dialectal gaps. Furthermore, this finding correlates with language script differences. We observe that 77.2% top-10 varieties in terms of maximum obtainable score are written with Latin script. Another finding is the performance instability of low-resource varieties across tasks. For instance, Old Guarani performs better in DEP parsing whereas, Mbyá Guarani (Paraguay) surpasses it in POS tagging even though the dataset remains the same (i.e. UD). For more detailed comparisons, in Appendix Table 21, we report the top-10 highest and lowest-scoring varieties across different tasks.
@@
Low-resource varieties written in Latin script receive greater benefit in zero-shot because of effective transfer from high-resource Standard English. With the presence of in-cluster/variety finetuning data, we effectively diminish this script effect to some extent. For example in the Hindi cluster, Fiji performs better than its Latin non-standard counterparts with in-cluster finetuning for NER (Table 12). In summary, if the standard variety of a cluster is non-latin but high-resource, then the success rate of in-cluster/variety fine-tuning tends to be higher. However, where all dialects are low-resource, Latin script varieties utilizing zero-shot transfer, eventually surpass others in the performance hierarchy. As an example, we report the zero-shot NER instances where the low-resource varieties perform better than the representative ones in Appendix Table 22 (most of these use Latin script).
@@ Appendix C Varieties and Clusters of DialectBench
    DialectBench Variety list (Table 8)
    Language Clusters and Representative Varieties (Table 9)
@@ Appendix D Result Visualizations
Dialectal Gap visualization for language clusters utilizing zero-shot cross-lingual transfer from Standard English.
@@ Appendix E Task Specific Evaluation Result Tables
    Named entity recognition (Table 12)
@@
    Machine translation (Table 20)

### v2

@@ 1 Introduction
Figure 2: DialectBench language clusters with their variety counts per task. "Other" encompasses 18 clusters (full cluster list in Appendix LABEL:lang_list).
@@ 2 DialectBench
Category Task Metric Source Dataset
Structured DEP parsing UAS Universal Dependency Zeman et al. (2021), TwitterAAE Blodgett et al. (2018), Singlish Wang et al. (2017)
Prediction POS tagging F1 Universal Dependency Zeman et al. (2021), Singlish Wang et al. (2017), Noisy Dialects Blaschke et al. (2023)
NER F1 Wikiann Pan et al. (2017); Rahimi et al. (2019), Norwegian NER Johansen (2019)
Classification DId F1 MADAR Bouamor et al. (2018), DMT Jauhiainen et al. (2019), Greek Sababa and Stassopoulou (2018), DSL-TL Zampieri et al. (2023), Swiss Germans Scherrer et al. (2019)
SA F1 TSAC Medhaffar et al. (2017), TUNIZI Fourati et al. (2021), DzSentiA Abdelli et al. (2019), SaudiBank Alqahtani et al. (2022), MAC Garouani and Kharroubi (2022), ASTD Nabil et al. (2015), AJGT Alomari et al. (2017), OCLAR Al Omari et al. (2019)
TC F1 SIB-200 Adelani et al. (2023)
NLI F1 XNLI Conneau et al. (2018) translate-test
Question MRC F1 Belebele Bandarkar et al. (2023)
Answering EQA Span F1 SDQA Faisal et al. (2021)
Generation MT BLEU CODET Alam et al. (2023), TIL-MT Mirzakhalov (2021)
@@ 3.2 Training and Evaluation
We use five general approaches for task-specific model training:
@@
    Task In-variety FT In-cluster FT Combined FT Zero-shot No ref-erence In-context learning
    DEP $\checkmark$ $\checkmark$
    POS $\checkmark$ $\checkmark$
    NER $\checkmark$ $\checkmark$
    EQA $\checkmark$ $\checkmark$ $\checkmark$
    MRC $\checkmark$
    NLI $\checkmark$
    TC $\checkmark$ $\checkmark$
    SA $\checkmark$ $\checkmark$
    DId $\checkmark$
    MT $\checkmark$ $\checkmark$
@@ 4 Results
Structured prediction DEP parsing 16 40 64.3 sw. shifted romance/brazilian portuguese 94.4 tupi-guarani sg./mbyá guaraní (a:brazil) 9.0
POS tagging 17 51 72.1 norwegian/norwegian bokmål (m:written) 98.7 tupi-guarani sg./mbyá guaraní (a:brazil) 1.9
@@
TC 15 38 77.7 sinitic/cmm. sinitic (o:traditional) 89.8 kurdish/central kurdish 19.4
DId 6 49 67.0 sinitic/mandarin chinese (a:taiwan, o:simp.) 98.6 sw. shifted romance/portuguese (m:written) 17.4
@@
Question Answering MRC 4 11 40.9 anglic/english 53.4 sotho-tswana (s.30)/southern sotho 29.0
@@ 4.1 Maximum Obtainable Scores
Here we provide key findings from our evaluation on each task. A task-specific summary is reported in Table 3. Detailed results comprising all tasks, models, language clusters and varieties are reported in Tables 10 to LABEL:zero-shot-mt in Appendix E.
@@
We present visualizations for the task-specific maximum scores. We show the one for Dependency Parsing in Fig. 3, where we observe that low-resource varieties from language clusters such as Tupi-Guarani (indigenous South American cluster), Saami and Komi (low-resource Uralic language clusters) have the lowest performance compared to Standard English and other closely related Germanic and Romance clusters. These low-resource varieties are also not included in the pretraining stage of our base language models (eg. mBERT). Furthermore, this trend is evident across all three structured prediction tasks. On the other hand, high-resource Indo-European languages such as Portuguese, French, and Norwegian usually perform better.
@@
For DId and SA, we generally collate different datasets for each language cluster and therefore, report the comparative classification results together. As a result, the locality level (e.g. city/region/country) also varies across clusters. For example, we report city-level DId results for Arabic and High German but country-level results for Portuguese, Spanish and English. In the case of SA, we have region/country-level results for Arabic varieties. For TC and NLI, we have the same set of clusters and varieties. However, we only report the zero-shot transfer performance from Standard English for NLI using the newly created translate-test NLI dataset.
@@
We generally do not see large gaps in performance within varieties in each language cluster. In EQA zero-shot experiments, English and its varieties have the highest performance overall and Korean varieties score the lowest. Combined fine-tuning boosts performance on all language clusters except in English. It’s important to note that these EQA scores primarily indicate the model’s robustness to accent-level differences and transcription noise, rather than broad dialectal robustness. However, further investigation is needed to determine whether this robustness specifically applies to both accent-level differences and transcription noise, or to any character-level variation up to a certain threshold.
For the MRC task, the performance peaked at 53.4 for Standard English, while the lowest score was 29 for Southern Sotho. More detailed results are presented in Tables 19 and 15.
@@ 4.2 Dialectal Gap Across Language Clusters
Task Gap metric Avg. val cluster (max) cluster (min)
@@ 5 Discussion
The highest-performing varieties are mostly standard high-resource languages and a few high-resource dialects (Norwegian dialects) whereas, the majority of the lowest-performing language variants are low-resourced varieties. This clear distinction of language varieties points towards the large existence of in-cluster dialectal gaps. Furthermore, this finding correlates with language script differences. We observe that 77.2% of top-10 varieties in terms of maximum obtainable score are written with Latin script. Another finding is the performance instability of low-resource varieties across tasks. For instance, Old Guarani performs better in DEP parsing whereas, Mbyá Guarani (Paraguay) surpasses it in POS tagging even though the dataset remains the same (i.e. UD). For more detailed comparisons, in Appendix Table 21, we report the top-10 highest and lowest-scoring varieties across different tasks.
@@
Low-resource varieties written in Latin script receive greater benefit in zero-shot because of effective transfer from high-resource Standard English. With the presence of In-cluster/variety finetuning data, we effectively diminish this script effect to some extent. For example in the Hindi cluster, Fiji performs better than its Latin non-standard counterparts with in-cluster finetuning for NER (LABEL:tab:ner_full). In summary, if the standard variety of a cluster is non-latin but high-resource, then the success rate of in-cluster/variety fine-tuning tends to be higher. However, where all dialects are low-resource, Latin script varieties utilizing zero-shot transfer, eventually surpass others in the performance hierarchy. As an example, we report the zero-shot NER instances where the low-resource varieties perform better than the representative ones in Appendix Table 22 (most of these use Latin script).
@@ 6 Conclusion
Acknowledgements
This material is based upon work supported by the US National Science Foundation under Grants No. IIS-2125466, IIS-2125948, CAREER Grant No. IIS2142739, as well as NSF Grants No. IIS-2203097 and IIS-2125201. We gratefully acknowledge support from Alfred P. Sloan Foundation Fellowship. This research is also supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200004. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix C Varieties and Clusters of DialectBench
    DialectBench Variety list (LABEL:lang_list)
    Language Clusters and Representative Varieties (LABEL:stdv)
@@ Appendix D Result Visualizations
Dialectal Gap visualization utilizing zero-shot cross-lingual transfer from Standard English.
@@ Appendix E Task Specific Evaluation Result Tables
    Named entity recognition (LABEL:tab:ner_full)
@@
    Machine translation (LABEL:zero-shot-mt)